OpenStudy (anonymous):

write an equation of the cosine function with aplitude of 2, period of 3pi, a phase shift of pi/6 and a midline at y=-4 (mostly need help with the period)

OpenStudy (anonymous):

*amplitude

OpenStudy (mertsj):

coefficient = 2pi/period

OpenStudy (dumbcow):

y = Acos(Bx+C) + k A- amplitude period = 2pi/B phase shift = Bx+C midline or vertical shift = k

OpenStudy (anonymous):

so how would i write the period?

OpenStudy (dumbcow):

just solve for the variables (coefficients) and plug them into the equation

OpenStudy (anonymous):

still confused..

OpenStudy (mertsj):

coefficient = 2pi/3pi=2/3

OpenStudy (anonymous):

oh ok thx

OpenStudy (mertsj):